# ASDF: Assembly State Detection Utilizing Late Fusion by Integrating 6D Pose Estimation

### v2

@@ Abstract
[…]
@@
Synthetic data, assembly state detection, 6D pose estimation.
[…]
- AR
@@ Introduction
Augmented Reality (AR) enables the dynamic visualization of object assembly guidance, presenting an innovative paradigm to support the assembly task execution. This can be realized through the utilization of instruction animations that seamlessly track 3D markers or leverage colliders to navigate through assembly steps [4, 8, 18]. An alternative approach to markers is the utilization of 6D pose estimation of assembly objects. These can combine assembly state detection and estimating the object poses to enable in-situ visualizations. By discerning assembly states, AR-driven in-situ visualization can provide guidance precisely tailored to the current assembly context.
@@
In order to provide in-situ guidance the assembly state is a relevant information and the knowledge of each objects position in space, namely its 6D object poses. Retrieving this knowledge is especially challenging in dynamic assembly scenarios [47, 37]. However, existing 6D pose estimation approaches [41, 25] are often limited to static scenes [16, 43] as the benchmark on which they are evaluated are static. Alternatively, object tracking is more often applied in dynamic scenarios with moving objects [35, 34, 19], but occlusion can be challenging for pure tracking-based approaches [19]. Dynamic scenarios with individual object assembly can occur in the medical or industry context [47, 18, 24]. There errors in assembly can lead to either wrong manufacturing, extended manufacturing times or in the medical context can even be challenging for ones life.
In that regard, AR guidance has tremendous potential. AR assembly guidance combined with 6D pose estimation and state-detection can enable in-situ guidance, and error detection during the assembly. Marker-based tracking, however, limits real-world applicability. To enable marker-less 6D pose estimation and state detection, using a convolutional neural network (CNN)-based approach can be beneficial. Existing 6D pose estimation approaches often focus on a single object [26, 41, 15, 35, 28] instead of assembly groups.
@@
Moreover, many existing deep learning-based assembly state detection approaches focus on 2D [31, 22, 48] or are limited in their evaluation to one or two objects [22, 24]. Moreover, current mulit-object 6D pose estimation approaches [41, 42] or tracking approaches [32, 35, 19] only provide the 6D object pose information.
0.1 Contribution
@@
To address these existing limitations we present Assembly State Detection utilizing late Fusion (ASDF) a deep learning-based assembly state and 6D pose estimation detection algorithm. We leverage deep learning building upon the real-time capable you only look once (YOLO) architecture fusing 6D pose estimation and assembly state detection for more precise object poses. For training we leverage purely synthetic data using online available 3D printing parts. We evaluate several aspects, such as network size and pose refinement on our ASDF dataset. Moreover, we compare our approach in 6D pose estimation performance on the Graph-based Object Tracking (GBOT) dataset. By evaluating the 6D pose estimation performance on this dataset, we can show first, the transfer ability from our domain-randomized training images to the GBOT evaluation images with different background, occlusion, blur and light conditions, and second, that the assembly state detection can lift the 6D pose performance.
@@
ASDF integrates late fusion in our Pose2State module, leading to refined assembly state detection and object pose estimation
A synthetic dataset encompassing 6D object poses and assembly states utilizing 3D printable part of reproducibility
Evaluation on two datasets, demonstrating the advantage of our approach compared to pure deep learning-based approaches
@@ 1 Related Work
Enabling in-situ guidance of assembly objects, requires the knowledge of an objects position in space, namely its 6D object pose, the knowledge of its assembly state and leveraging assembly data for training and evaluating deep learning-based approaches.
@@
Early approaches for instance-level 6D pose estimation applied template matching [11]. With the progress of deep learning based approaches (CNN, Transformer) one-stager [10, 46] and two-stager [41] approaches have been developed.
One-stager approaches are end-to-end trainable [10, 46] They extract features from a segmentation or object detection backbone. These can be regressed directly [46] or other output like keypoints can be feed to Perspective-n-Point (PnP) [38]/Least Squares Fitting [10]. Amini et al. [2] introduce YoloPose which directly regresses keypoints in an image and presents a learnable module to replace PnP.
Two-stager approaches, apply a state-of-the-art object detection algorithm, for example Faster R-CNN and build the 6D pose estimation on top of these predictions. Wang et al. [41] utilize geometric feature regression on top of the object detection algorithm. This results in 2D-3D correspondences and Surface Region Attention which is leveraged in Patch-PnP. Similarly, Pix2Pose [25] leverages 2D bounding boxes followed by mask prediction and bounding box refinement. The final result is predicted using RANSAC PnP.
In addition to deep learning-based 6D pose estimation, another field is object tracking. Object tracking starts with an initialization and then assumes the 6D pose in ongoing frames. Stoiber et al. [35] proposes Iterative Correspondence Geometry (ICG) combines visual, regions and depth information. Their improvement ICG [32] additionally considers SIFT and ORB features. Additionally, Mb-ICG [34] tackles kinematic structures. Object tracking can get lost in highly occluded scenes. There, combining deep learning-based approaches with tracking can be promising promising [20, 19]. For example, DeepIM [20] combines the re-initialization by PoseCNN [46]. Li et al. [19] propose a combination of an extension of YOLOv8 and ICG. While their pose initialization relies on the CNN, they utilize an assembly state-graph for tracking assembled objects. Additionally, they provide the GBOT dataset for tracking assembled parts. However, the focus lies on 6D pose estimation instead of assembly detection.
@@ 1.2 Assembly State Detection for Augmented Reality
While the object pose is on essential aspect for assembly guidance another on is the assembly state itself. Zauner et al. [47] utilized markers and build an assembly graph for AR-based assembly instruction. Duplotrack [9] utilizes point cloud alignment whenever the assembly building blocks change.
In terms of 2D assembly state detection, Kleinbeck et al. [18] combine YOLOv5 and synthetic data for building block assembly guidance. The guidance steps are rendered in a Hololens. Similarly, Stane et al. [31] provide AR glass based guidance. For their state detection they introduce a convolution block in YOLO. Overall, their state detection improved the YOLO-based object prediction.
Liu et al. [22] integrate a two-fold attention mechanism in a 2D object detection architecture. Zhou et al. [48] address mobile AR guidance utilizing regions-of-interest for state identification. The method is two-fold, in a first step the regions-of-interest are extracted, in a second step, the state recognition is trained.
Other approaches considering 3D/6D build upon the relative pose between objects [45, 23, 24]. Wu et al. [45] follow a tree like graph structure to determine the assembly state. Similarly, Murray et al. [24] address 6D pose estimation and assembly prediction in industrial settings defining an assembly state graph. Their network build upon Pix2Pose [25]. To address multi-view input they utilized depth estimation and project pixels to 3D. Su et al. [37] utilize a TridentNet as backbone and add a pose prediction head. They predict the pose and assembly state for one object with five assembly states and provide AR guidance for utilizing the network.
@@ 1.3 Real-world and Synthetic Datasets
For 6D pose estimation/object tracking multiple benchmarks exist, for example, one common benchmark is the YCB-V benchmark [46]. However, this and other common ones considers single objects without state changes [5, 46, 16, 14]. Moreover, providing 6D pose datasets with complex scenes is challenging. To address this one common approach can the utilization of synthetic data [29]. In the 6D pose domain one common approach for synthetic data generation [7, 27, 3, 29].
@@ 1.3.1 Assembly State Datasets
One common area for assembly datasets is IKEA furniture. The IKEA-Manual dataset [44] provides assembly data including the 3D pose of each part, specifically detailing the rotation of the 3D components. However, it does not explore the translation of individual parts. Addressing this limitation, the IKEA assembly dataset introduced by Su et al. [36] fills the gap, although it is not publicly accessible.
Other works build upon individual objects [24, 37]. However, this makes reproduciblity more challenging. Schoonbeek et al. [30] present a hybrid dataset focusing on action recognition and assembly state detection. They use multiple modalities (RGB, Depth, Gaze, Stereo) and record their dataset with a Hololens 2. Su et al. [37] reconstruct a real-world coffee machine and sample the parts on images to generate a synthetic assembly state dataset. The dataset features five states for the coffee machine assembly. Li et al. [19] propose the use of 3D printable objects and present a synthetic dataset featuring these objects. The ground truth contains the 6D object poses at several assembly states. However, the dataset focuses purely on the 6D pose.
@@ 1.3.2 Synthetic Data
The use of synthetic data for either pre-training a 6D pose estimation or purely training on synthetic data can be usefully as recording a real-world marker-less dataset with only a low annotation error can be challenging [16, 29]. For generating synthetic data in the context of 6D pose estimation/assembly state detection or in general object detection one challenge is the so-called sim-to-real gap [29, 39]. This gap denotes the domain gap between synthetic images and real world captures. To address this gap one common approach is randomizing several parameters like light, background and object textures [29, 39]. Moreover, to focus on the objects of interest but not purely over-fit on them adding distracting objects to the training set can be beneficial [29, 40, 1]. Tremblay et al. [40] investigated this by combining realistic and randomized images and added geometric distracting objects leading to, i. e. occlusion of the actual object of interest. This lifted the object detection performance. Alghonhaim et al. [1] benchmarked domain randomization considering background, textures and distractors. They proved as well that distractors are benefical for generalization.
While the constant progress in 6D pose estimation is promising for deep learning-based approaches for single objects, the combination of 6D pose estimation and assembly state detection is even more promising for AR guidance approaches. However, for this tasks the used objects hugely vary from building blocks up to engines with over a hundred assembly states. Therefore, we contribute a comparable approach to existing 6D pose estimation/object tracking approaches and a dataset building upon these re-usable assembly assets to improve comparability in this domain.
@@ 2.1 Assembly State Detection utilizing late Fusion
ASDF combines 6D pose estimation and assembly state detection, building upon YOLOv8Pose [19]. RGB-D images serve as input, with the backbone processing the pure RGB image while leveraging the depth image and resulting point cloud for pose refinement. The system encompasses both 6D pose estimation and assembly state detection, which we elaborate on separately. These distinct components are fused within our Pose2State module, as depicted in Figure 1.
@@
For 6D pose estimation, we utilize YOLOv8Pose [19], which extends the output of YOLOv8 using RANSAC PnP. To establish our keypoints in 3D space, we utilize the 2D keypoints predicted by the backbone. Employing Farthest Point Sampling (FPS) [19, 26], we distribute the points as widely apart from each other as possible. For the final number of keypoints (17), we strike a balance between computational cost (as more keypoints lead to higher computation cost) and performance.
@@
Our assembly pose refinement step further refines the estimated translation by determining the boundary values of each assembly in each assembly group, as illustrated in Figure 1. To achieve this, we determine the required movement along one coordinate axis (z-axis) and scale the other two axes accordingly. For calculating this movement perpendicular to the camera plane, we transform the 3D surface points of the component using the transformation matrix obtained from the initial pose estimation by our CNN ().
The transformed points of the 3D surface points using can be expressed as:
@@
Finally, based on the selected points in and the approximated depth, we determine the required movement along one coordinate axis (i. e., ) and scale the other two axes (i. e., and ) accordingly to refine the translation part. Using the approximated 3D points, we calculate the difference to the corresponding 3D surface points. This allows estimating the shift for all points using:
denotes the estimated shift distance for point and represents the total number of points. is the the overall estimate of the required shift for all points.
To estimate the shift distance for each point, we use . Then, to obtain an overall estimate, we combine these individual estimations using a weighting function . This function assigns higher weights to points with smaller differences. This weighting prohibits that occluded objects retrieve a false calculation during the refinement step.
The resulting translation along the axis perpendicular to the image plane can now be used to determine the vector along the camera view axis that represents the final translation.
@@ 2.1.2 Assembly State Detection
The final assembly state detection combines three key components: deep learning-based state detection, relative pose-based state detection, and consideration of the previous state, employing a weighted combination approach.
@@
The assembly state detection system utilizes both relative pose information and deep learning-driven state detection. To seamlessly integrate state detection into our framework, we represent each state as a distinct class for the network to learn. Additionally, to enhance prediction accuracy, we maintain a assembly state recording for the whole assembly. For each part of the assembly we keep a record if pose matrix exists.
@@
The relative pose between objects offers valuable information regarding the current state of each assembly. These relative poses are inherently known, as they are crucial for generating ground truth training and evaluation data.
To determine the assembly state based on the relative pose from one assembly, we adopt a strategy of selecting the base part of each assembly. This base part serves as a reference point from which the relative poses of the other parts are considered. By anchoring the analysis to these base parts, we can effectively interpret the relative poses of the entire assembly, allowing for accurate assessment and monitoring of its overall configuration and state.

### v3

@@ Abstract
In medical and industrial domains, providing guidance for assembly processes can be critical to ensure efficiency and safety. Errors in assembly can lead to significant consequences such as extended surgery times and prolonged manufacturing or maintenance times in industry. Assembly scenarios can benefit from in-situ augmented reality visualization, i.e., augmentations in close proximity to the target object, to provide guidance, reduce assembly times, and minimize errors. In order to enable in-situ visualization, 6D pose estimation can be leveraged to identify the correct location for an augmentation. Existing 6D pose estimation techniques primarily focus on individual objects and static captures. However, assembly scenarios have various dynamics, including occlusion during assembly and dynamics in the appearance of assembly objects. Existing work focus either on object detection combined with state detection, or focus purely on the pose estimation. To address the challenges of 6D pose estimation in combination with assembly state detection, our approach ASDF builds upon the strengths of YOLOv8, a real-time capable object detection framework. We extend this framework, refine the object pose, and fuse pose knowledge with network-detected pose information. Utilizing our late fusion in our Pose2State module results in refined 6D pose estimation and assembly state detection. By combining both pose and state information, our Pose2State module predicts the final assembly state with precision. The evaluation of our ASDF dataset shows that our Pose2State module leads to an improved assembly state detection and that the improvement of the assembly state further leads to a more robust 6D pose estimation. Moreover, on the GBOT dataset, we outperform the pure deep learning-based network and even outperform the hybrid and pure tracking-based approaches.
@@
6D pose estimation, assembly state detection, synthetic data- AR
@@ Introduction
Providing error-free assembly of complex assembly groups is highly relevant in manufacturing, maintenance, or medical scenarios [4, 51, 19, 25]. Object assembly in these cases can be challenging due to the time pressure, the need for high accuracy, and the requirement for prior knowledge about the individual assembly parts. Moreover, assembly errors in manufacturing can lead to broken parts, extended manufacturing times or in the medical context this can affect the surgery time. Support during assembly tasks can reduce physiological and psychological loads in such scenarios [39, 11]. To support these scenarios, augmented reality (AR) guidance can provide valuable dynamic visualization during the assembly process. Enabling dynamic visualization for assembly processes requires accurate real-time tracking of the objects and their current state in the assembly. While markers proved high accuracy for tracking in AR they are sometimes not suitable for every task in manufacturing or medical scenarios [4]. An alternative approach to markers in AR, is the utilization of 6D pose estimation of assembly objects [21]. Deep learning-driven 6D pose estimation determines both the position and orientation of objects in three-dimensional space, enabling markerless tracking.
@@
To provide in-situ guidance, knowledge of the current assembly state is required along with each objects position in space. Retrieving this knowledge is especially challenging in dynamic assembly scenarios [51, 38]. Furthermore, existing 6D pose estimation approaches [45, 26] are often limited to static scenes [17, 46] as the benchmark on which they are evaluated are static. Alternatively, object tracking is more often applied in dynamic scenarios with moving objects [36, 35, 21], but occlusion can be challenging for pure tracking-based approaches [21]. AR assembly guidance combined with 6D pose estimation and state-detection can enable in-situ guidance, and error detection during the assembly to reduce risks and assembly time. Marker-based tracking, however, limits real-world applicability. To enable marker-less 6D pose estimation and state detection, using a deep learning-based approach can be beneficial.
@@
In summary, existing deep learning-based multi-object approaches a) limit their object and state detection to 2D [23, 32, 53], b) limit their evaluation to one or two objects [23, 25], or c) only provide the 6D object pose information without predicting the current assembly state [21, 33, 36].
@@
To address these limitations, we present assembly state detection utilizing late fusion (ASDF), a deep learning-based system for assembly state detection and 6D pose estimation. We build upon the real-time capable you only look once (YOLO) architecture, fusing 6D pose estimation and assembly state detection for more precise object poses. Additionally, we propose a fully synthetic dataset for training and evaluation, with an additional real-world test scene. Our ASDF dataset consists of online available 3D printing parts with 6D pose estimation ground truth and assembly state ground truth. The test dataset contains full assembly sequences with hand occlusion and faulty states for a robust evaluation. We evaluate various aspects, such as network size and pose refinement, on our ASDF dataset.
Furthermore, we compare our approach’s 6D pose estimation performance on the graph-based object tracking (GBOT) dataset [21]. This evaluation demonstrates the transferability of our domain-randomized training images to the GBOT evaluation images and shows that assembly state detection can enhance 6D pose performance.
@@
ASDF, a late fusion approach enhancing assembly state detection and 6D pose estimation through improved assembly state prediction
Our synthetic ASDF dataset that includes 6D object poses and assembly states using 3D printable parts for reproducibility111GitHub ASDF
State-of-the-art results on two datasets with assembly assets, demonstrating the advantage of our approach
@@ 1 Related Work
Instance-level 6D pose estimation can be divided in one-stager [10, 49] and two-stager [45] approaches. One-stagers are end-to-end trainable [10, 49]. They extract features from a segmentation or object detection backbone. These can be regressed directly [49] or other output like keypoints can be feed to perspective-n-point (PnP) [40]/Least Squares Fitting [10]. Amini et al. [2] introduce YoloPose which directly regresses keypoints in an image and presents a learnable module to replace PnP. Two-stagers, apply a state-of-the-art object detection algorithm, for example Faster R-CNN [8] and build the 6D pose estimation on top of these predictions. Wang et al. [45] utilize geometric feature regression on top of the object detection algorithm. This results in 2D-3D correspondences and Surface Region Attention which is leveraged in Patch-PnP. Similarly, Pix2Pose [26] leverages 2D bounding boxes followed by mask prediction step and bounding box refinement step. The final result is predicted using RANSAC PnP [7, 20, 52]. While one-stagers are often more computationally cheap during inference time, two-stagers can provide more accuracy.
In addition to deep learning-based 6D pose estimation, another crucial field for AR-driven instructions and object-based applications is object tracking [44]. Object tracking starts with a pose initialization and then assumes the 6D object pose in the following frames. Stoiber et al. [36] combine visual, regions and depth information. Their improvement ICG [33] additionally considers SIFT and ORB features. For objects consisting of multiple parts, Mb-ICG [35] looked at kinematic structures. However, they track the complete assembled object instead of the individual assembly steps. Object tracking can get lost in highly occluded scenes. Li et al. [21] propose a combination of an extension of YOLOv8 and an improvement of iterative correspondence geometry (ICG) to re-initialize the tracking if it gets lost. While their pose initialization relies on the convolutional neural network (CNN), they utilize an assembly state-graph for tracking assembled objects. Furthermore, they provide the GBOT dataset for tracking assembled parts.
@@ 1.2 Assembly State Detection for Augmented Reality
While the object pose is on essential aspect for assembly guidance another on is the assembly state itself. Zauner et al. [51] utilized markers and build an assembly graph for AR-based assembly instruction. Duplotrack [9] utilizes point cloud alignment whenever the assembly building blocks change. Radkowski et al. [28], utilize object tracking with iterative closest point (ICP)-based refinment for AR-guided assembly matching point clouds.
In terms of 2D assembly state detection, Kleinbeck et al. [19] combine YOLOv5 [16] and synthetic data for building block assembly guidance. The guidance steps are rendered in a Hololens. Similarly, Stanescu et al. [32] provide AR glasses-based guidance. For their state detection they introduce an extra convolution block in YOLO. Overall, their state detection improved the YOLO-based object prediction.
Liu et al. [23] integrate a two-fold attention mechanism in a 2D object detection architecture.They test their approach on two objects, an IKEA table assembly and Fender assembly. Zhou et al. [53] address mobile AR guidance utilizing regions-of-interest for state identification. The method is two-fold, in a first step the regions-of-interest are extracted, in a second step, the state recognition is trained.
Other approaches considering 3D/6D build upon the relative pose between objects [24, 25, 48]. Wu et al. [48] follow a tree like graph structure to determine the assembly state. Similarly, Murray et al. [25] address 6D pose estimation and assembly prediction for robot bin picking defining an assembly state graph. They omit training object detection as a first stage since they are dealing with one single object in front of the camera. For 6D pose estimation they follow Pix2Pose [26] and address multi-view input utilizing depth estimation and project pixels to 3D. Su et al. [38] utilize a TridentNet as backbone and add a pose prediction head. They predict the pose and assembly state for one object with five assembly states and provide AR guidance for utilizing the network.
@@ 1.3 Real-world and Synthetic Datasets
For 6D pose estimation/object tracking multiple benchmarks exist, for example, one common benchmark is the YCB-V benchmark [49]. However, this and other common ones considers single objects without state changes [3, 49, 17, 15]. Moreover, providing 6D pose datasets with complex scenes and switching object states is challenging. To address this challenge, one common approach is the use of synthetic data [29, 25, 21].
@@ 1.3.1 Assembly State Datasets
One common area for assembly datasets is IKEA furniture [47, 37, 23]. The IKEA-Manual dataset [47] provides assembly data including the 3D pose of each part, specifically detailing the rotation of the 3D components. However, it does not explore the translation of individual parts. Addressing this limitation, the IKEA assembly dataset [37] fills the gap, although it is not publicly accessible.
Other works build upon individual objects [25, 38, 32]. However, this makes reproduciblity more challenging. Schoonbeek et al. [30] present a hybrid multi-modal dataset focusing on action recognition recorded with a Hololens 2. Su et al. [38] reconstruct a real-world coffee machine and sample the parts on images to generate a synthetic assembly state dataset.
For reproducibility, Li et al. [21] propose the use of 3D printable objects and present a synthetic dataset featuring these objects and their poses. The individual objects have a varying number of assembly states and different sizes.
@@ 1.3.2 Synthetic Data
The use of synthetic data for 6D position estimation can be useful as obtaining a real markerless dataset is a challenging task [17, 29, 41]. The use of synthetic data poses the challenge of the so-called sim-to-real gap [29, 42]. This gap refers to the domain gap between synthetic and real images. To close this gap, various parameters such as light, background, object textures are often randomized and distracting objects are added [29, 43, 1]. Tremblay et al. [43] investigated this by combining realistic and randomized images and adding distractor objects. This increased the performance of object recognition. Alghonhaim et al. [1] tested domain randomization considering background, textures and distractors. They also proved that distractors are beneficial for the generalization of a CNN.
The constant progress in 6D pose estimation is promising for single objects [45, 40, 27, 26, 10, 50], the combination of 6D pose estimation and assembly state detection is even more promising for AR guidance approaches [21, 25]. However, for this tasks the used objects hugely vary from building blocks [39, 19, 32] up to engines with over a hundred assembly states [25]. Existing approaches are often limited to 2D [32], omit training an object detection model in the first stage [25], require additional state recognition in addition to their robust object tracking [21] or can not handle state switches [35]. To address these limitations we propose ASDF an end-to-end approach for assembly state detection and 6D pose estimation. Moreover, existing datasets for this task are again either for 2D object detection, 6D pose estimation [21], limited to action recognition [31] or not publicly available. To enhance comparison in this area we propose our ASDF dataset containing 6D object pose ground and assembly state truth data using 3D printable object’s for comparability and reproducibility.
@@ 2.1 Assembly State Detection Utilizing Late Fusion
ASDF estimates both the 6D pose and the assembly state, which we will explain separately. The individual outputs are fused in a late fusion step within our Pose2State module.
@@
ASDF uses RGB-D images. The backbone processes the pure RGB image. The Translation Refinement module utilizes the depth image and resulting point cloud to refine the pose, see Figure 1.
For 6D pose estimation, we leverage YOLOv8Pose [21], an extension of YOLOv8 utilizing RANSAC PnP. To establish keypoints in 3D space, we utilize the 2D keypoints predicted by the backbone. Using farthest point sampling (FPS) [21, 27], we distribute the points as widely apart as possible. The selection of a final number of keypoints (17) creates a balance between computing costs and performance, as more keypoints increase the computing costs.
@@
Our assembly pose refinement step enhances the estimated translation by identifying the boundary values of each assembly in every assembly group, see Figure 1. To accomplish this, we determine the necessary movement along one coordinate axis (the z-axis) and proportionally adjust the other two axes. To calculate this movement perpendicular to the camera plane, we transform the 3D surface points of the component using the transformation matrix obtained from the initial pose estimation by our CNN ().
The transformed points of the 3D surface points using can be expressed as follows:
@@
Finally, based on the selected points in and the approximated depth, we determine the necessary movement along one coordinate axis (i.e., ) and proportionally adjust the other two axes (i.e., and ) to refine the translation part. By utilizing the approximated 3D points, we calculate the difference from the corresponding 3D surface points. This allows us to estimate the shift for all points using:
denotes the estimated shift distance for point and represents the total number of points. is the overall estimate of the required shift for all points.
To estimate the shift distance for each point, we use . Then, to obtain an overall estimate, we combine these individual estimations using a weighting function . This function assigns higher weights to points with smaller differences, thereby preventing occluded objects from influencing false calculations during the refinement step.
The resulting translation along the axis perpendicular to the image plane can now be utilized to determine the vector along the camera view axis, representing the final translation.
@@ 2.1.2 Assembly State Detection
The final assembly state detection combines three key components: deep learning-based state detection, relative pose-based state detection, and consideration of the previous state using a weighted late fusion.
@@
To integrate state detection in ASDFs backbone, we represent each state as a distinct class. The assembled parts with a new assembly state are regarded as a new class.
@@
The relative position between objects provides valuable information about the current assembly state of each assembly. These relative poses are known because they are crucial for the creation of training and evaluation data.
To determine the assembly state based on the relative pose of an assembly, we select the base part of each assembly. Followed by this, we can interpret the relative poses of the entire assembly.